The mass of a certain plane (p) is 1/100 that of the earth, while the radius is 1/4 that of the earth's radius. If the acceleration due to gravity of the earth is 10m/s^2, what is the value of p.
1
Expert's answer
2019-09-27T09:58:19-0400

Acceleration ratio "\frac{a_p}{g} = (\frac{R}{r_p})^2(\frac{m_p}{M})"


which means "a_p = 10\cdot16\cdot10^{-2} = 1.6 (m/s^2)"